# Multicell-Fold: geometric learning in folding multicellular life

### v1

@@ Abstract
During developmental processes such as embryogenesis, how a group of cells fold into specific structures, is a central question in biology that defines how living organisms form [1, 2, 3, 4, 5]. Establishing tissue-level morphology critically relies on how every single cell decides to position itself relative to its neighboring cells [6, 7]. Despite its importance, it remains a major challenge to understand and predict the behavior of every cell within the living tissue over time during such intricate processes. To tackle this question, we propose a geometric deep learning model that can predict multicellular folding and embryogenesis, accurately capturing the highly convoluted spatial interactions among cells. We demonstrate that multicellular data can be represented with both granular and foam-like physical pictures through a unified graph data structure, considering both cellular interactions and cell junction networks. We successfully use our model to achieve two important tasks, interpretable 4-D morphological sequence alignment, and predicting local cell rearrangements before they occur at single-cell resolution. Furthermore, using an activation map and ablation studies, we demonstrate that cell geometries and cell junction networks together regulate local cell rearrangement which is critical for embryo morphogenesis. This approach provides a novel paradigm to study morphogenesis, highlighting a unified data structure and harnessing the power of geometric deep learning to accurately model the mechanisms and behaviors of cells during development. It offers a pathway toward creating a unified dynamic morphological atlas for a variety of developmental processes such as embryogenesis.
Two-sentence summary: We present a foundational data structure and geometric deep-learning workflow for deciphering the folding of multicellular life. In 3-D full-embryo video data, we demonstrate geometric video sequence alignment, and highly accurate prediction of cell rearrangements prior to their occurrence at single-cell resolution.
@@
Here we present a foundational geometric deep-learning workflow tailored to predict mesoscale multicellular evolution, where both the granular and foam-like physical pictures can be represented within a unified dual-graph data structure (Fig. 1). With this modeling approach, for a time-lapsed Drosophila 3-D full-embryo dataset, we first demonstrate unsupervised and interpretable geometric video sequence alignment (Fig. 2), which enables large-scale 4-D morphological sequencing and data registration. Furthermore, a model is trained on one 3-D Drosophila whole-embryo video with labeled cell rearrangements, and then validated on a different Drosophila whole-embryo video to identify future rearrangements that have not yet happened, and we demonstrate an unprecedented ability to predict cell rearrangements at single-cell resolution before they occur, with an overall accuracy $>82\%$ (Fig. 3, Supplementary Materials Table S1 and Table S2). Our workflow is the first algorithm, to our knowledge, that can achieve close-to-experiment accuracy in predicting multicellular folding processes at single-cell resolution. Beyond prediction, using an analysis of neural activation maps and ablation studies, we also identify key geometric signatures underlying these crucial developmental processes.
@@
To demonstrate this, we perform video sequence alignment of two 3-D time-lapsed whole Drosophila embryos during gastrulation from a recent experimental study [8]; two embryos are imaged at arbitrary orientations and different frame rates, and they start to evolve at different times (Supplementary Materials Fig. S4). Using our workflow, we train a regressional model on one embryo as a reference, with its own frame ID as labels. Applying the model to each frame of the second embryo outputs the ID of the most aligned frame from the first embryo, which aligns the developmental time of the two embryos (Fig. 2a). To interpret the alignment results, in a trained model, we adapt the class activation method [52]; we visualize the weighed activation of the last graph encoding layer before the global pooling operation. The activation map shows clearly that the model captures the ventral furrow and invaginated apical and dorsal features (Fig. 2b). An impressive fact is that the model has learned to focus on these features even though it is only trained to regress frame ID, which is very little information.
@@
Dataset.— We use a publicly available Drosophila gastrulation dataset published in [8]. The raw dataset contains segmented time-lapsed 3-D full embryo configurations throughout the gastrulation process. Specifically, for each frame, the dataset contains cell coordinates, cell vertex coordinates, cell-cell adjacency, and cell junctions. The data is pre-processed using a customized MATLAB script to arrive at the input format for our geometric deep-learning model. The script for pre-processing the data can be found on our GitHub site.
@@ Acknowledgements
Our code will be publicly available on our GitHub repository at: https://github.com/GuoLab-CellMechanics/Multicell-Fold. Trained model weights will be available on Hugging Face.

### v2

@@ Abstract
During developmental processes such as embryogenesis, how a group of cells fold into specific structures, is a central question in biology. However, it remains a major challenge to understand and predict the behavior of every cell within the living tissue over time during such intricate processes. Here we present a geometric deep-learning model that can accurately capture the highly convoluted interactions among cells. We demonstrate that multicellular data can be represented with both granular and foam-like physical pictures through a unified graph data structure, considering both cellular interactions and cell junction networks. Using this model, we achieve interpretable 4-D morphological sequence alignment, and predicting cell rearrangements before they occur at single-cell resolution. Furthermore, using neural activation map and ablation studies, we demonstrate cell geometries and cell junction networks together regulate morphogenesis at single-cell precision. This approach offers a pathway toward a unified dynamic atlas for a variety of developmental processes.
@@
Here we present a foundational geometric deep-learning workflow tailored to predict mesoscale multicellular evolution, where both the granular and foam-like physical pictures can be represented within a unified dual-graph data structure (Fig. 1). With this modeling approach, for a time-lapsed Drosophila 3-D whole-embryo dataset, we first demonstrate unsupervised and interpretable geometric video sequence alignment (Fig. 2), which enables large-scale 4-D morphological sequencing and data registration. Furthermore, a model is trained on one 3-D Drosophila whole-embryo video with labeled cell rearrangements, and then validated on a different Drosophila whole-embryo video to identify future rearrangements that have not yet happened, and we demonstrate an unprecedented ability to predict cell rearrangements at single-cell resolution before they occur, with an overall accuracy $>82\%$ (Fig. 3, Supplementary Materials Table S1 and Table S2). Our workflow is the first algorithm, to our knowledge, that can achieve close-to-experiment accuracy in predicting multicellular folding processes at single-cell resolution. Beyond prediction, using an analysis of neural activation maps and ablation studies, we also identify key geometric signatures underlying these crucial developmental processes.
@@
To demonstrate this, we perform video sequence alignment of two 3-D time-lapsed whole Drosophila embryos during gastrulation from a recent experimental study [7]; two embryos are imaged at arbitrary orientations and different frame rates, and they start to evolve at different times (Supplementary Materials Fig. S4). Using our workflow, we train a regressional model on one embryo as a reference, with its own frame ID as labels. Applying the model to each frame of the second embryo outputs the ID of the most aligned frame from the first embryo, which aligns the developmental time of the two embryos (Fig. 2a). To interpret the alignment results, in a trained model, we adapt the class activation method [52]; we visualize the weighed activation of the last graph encoding layer before the global pooling operation. The activation map shows clearly that the model captures the ventral furrow and dorsal features (Fig. 2b). An impressive fact is that the model has learned to focus on these features even though it is only trained to regress frame ID, which is very little information.
@@
Dataset.— We use a publicly available Drosophila gastrulation dataset published in [7]. The raw dataset contains segmented time-lapsed 3-D whole embryo configurations throughout the gastrulation process. Specifically, for each frame, the dataset contains cell coordinates, cell vertex coordinates, cell-cell adjacency, and cell junctions. The data is pre-processed using a customized MATLAB script to arrive at the input format for our geometric deep-learning model. The script for pre-processing the data can be found on our GitHub site.
@@ Acknowledgements
Our code will be publicly available on our GitHub repository at:
Trained model weights will be available on Hugging Face.